# Natural Fingerprints of Large Language Models

### v1

@@ 4.5 SFT experiment
5 Discussion and Conclusion
In this work, we explored the origins of bias in LLMs by studying whether their outputs contain model-specific patterns, which we refer to as natural fingerprints.
Our experiments confirmed that different publicly available pretrained language models exhibit detectable biases in their outputs. This bias appears to stem largely from differences in the training data. Since LLMs are generative models trained to imitate the distribution of their training data, the results are intuitive.
However, we also found that even when models are trained on the same dataset, they still produce distinguishable outputs. This surprising result suggests that other factors contribute to these differences. Through controlled experiments, we showed that certain hyperparameters, such as learning rate, weight decay, and initial weights, as well as the order of the training data, are enough to introduce detectable biases in the model outputs. The sensitivity of LLMs to training data order aligns with recent training practices. Some models are trained using a multi-stage approach, where higher-quality data is presented later in the training process to improve overall performance. This approach assumes that the model becomes more influenced by the data it sees in the later stages of training.
We believe that a deeper understanding of these contributing factors is essential for designing more capable and fair LLMs in the future.
Appendix A Pretraining Hyperparameters
Table 6 shows that hyperparameters used in training our LLMs from scratch. We use SwiGLU (Shazeer, 2020) as the activation function of FFN layers, and apply grouped query attention Ainslie et al. (2023) to self-attention layers. We use scaled embed (Takase et al., 2024) to train some models with high learning rate: 5e-4.
@@ Appendix A Pretraining Hyperparameters
$lr$ decay style cosine
$lr$ warmup step 2500
@@ Appendix B Fine-tuning Hyperparameters
$lr$ decay style cosine
$lr$ Warmup ratio 0.002
$lr$ decay ratio 0.9
@@
Appendix C Downstream Performance of the Trained Models
Hyperparameter
lr wd seed PIQA HellaSwag WinoGrande ARC-e ARC-c OBQA
5e-4 0.1 1 71.6 52.2 54.1 63.4 35.0 37.6
5e-4 0.1 2 70.8 51.4 54.1 61.7 34.6 36.0
5e-4 0.1 3 71.0 52.0 54.3 62.2 34.0 39.6
1e-3 0.1 1 71.5 53.8 55.5 64.8 36.4 38.4
1e-4 0.1 1 69.9 46.4 53.9 59.7 31.8 35.6
5e-4 0.01 1 70.4 50.5 52.2 60.7 33.9 37.4
5e-4 0 1 70.7 51.2 54.4 62.2 34.2 37.8
In Section 4.3, we trained multiple LLMs with different hyperparameters. Table 8 shows the zero-shot performance of these LLMs on common sense reasoning tasks: PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019a), WinoGrande (Sakaguchi et al., 2021), ARC easy and challenge (Clark et al., 2018), and OpenBookQA (OBQA) Mihaylov et al. (2018). The model trained with the lowest learning rate (1e-4) achieves the worst performance across all tasks except PIQA, while the other LLMs perform comparably. Overall, the trained LLMs demonstrate similar quality on downstream tasks.

### v2

@@ 4.5 Do Natural Fingerprints Emerge in Supervised Fine-Tuning?
Setting Model Unigram feature
Model1 uis _admit _== _elif ”. _pause _gentleman : eh script
Order Model2 perate ah _folder D 1 _ac 8 _Popular Focus vis
Model3 _Before _exclus _case _saw _Adam ]) ; _bell _Mike ();
Model1 _Tommy ceil _the scar apple HTTP _Sam Over Self In
Init Model2 __ [ _Forces _result urchase !. ovi _YES headers 0
Model3 One : _– Fin ’ This _Louisiana ()) _sep _}
Table 7: The top 10 largest weighted features to detect each fine-tuned LLM. We represent a whitespace with _.
@@
Table 7 shows the top 10 largest weighted unigram features used to detect each fine-tuned LLM. Notably, many of the top unigram features appear uncommon in ordinary conversational contexts, which may reflect idiosyncrasies of the Alpaca dataset used for SFT. Since Alpaca consists of instructions and demonstrations generated via the text-davinci-003 engine with aggressive decoding and a simplified pipeline, some tokens may originate from atypical or templated instruction formats rather than natural dialogue. This suggests that part of the distinguishability among models may stem from how these non-natural tokens are memorized or reproduced differently depending on initialization or training dynamics.
5 Conclusion
In revealing that training setups alone can induce distinguishable patterns in LLM outputs, even when data and architecture are held constant, our study highlights a previously underexplored source of variability in generative models. This phenomenon, termed natural fingerprints, highlights subtle yet consistent differences arising from minor training variations, including parameter initialization and training data order, even under identical datasets and architectures.
This suggests that future research on transparency, reliability, and interpretability should also consider training dynamics, not just dataset differences, as a source of model-specific behavior. We hope that this work serves as a stepping stone toward a deeper understanding of the structural signals embedded in LLM outputs—and that it invites further exploration into how models subtly encode the traces of how they were trained.
Appendix: Natural Fingerprints of Large Language Models
Appendix A Hyperparameters
Table 8 shows that hyperparameters used in training our LLMs from scratch. We use SwiGLU (Shazeer 2020) as the activation function of FFN layers, and apply grouped query attention (Ainslie, Lee-Thorp et al. 2023) to self-attention layers. We use scaled embed (Takase, Kiyono et al. 2024) to train some models with high learning rate: 5e-4.
@@ Appendix A Hyperparameters
lrlr decay style cosine
lrlr warmup step 2500
@@
lrlr decay style cosine
lrlr Warmup ratio 0.002
lrlr decay ratio 0.9